OpenStudy (anonymous):

HELP ME Solve the equation for x. Do not include "x = " in your answer. If necessary, enter a non-integer result as a fraction using the slash ( / ) bar. The square root 2-3x divided by the square root 5x =3

OpenStudy (paxpolaris):

The square root 2-3x divided by the square root 5x =3 \[\Large {\sqrt{2-3x} \over \sqrt{5x}}=3\]

OpenStudy (paxpolaris):

\[\Large \implies{\sqrt{2-3x} }=3\sqrt{5x}\] then square both sides

OpenStudy (anonymous):